OpenStudy (anonymous):

Find four consecutive integers such that the sum of the second and the fourth is 132.

OpenStudy (2bornot2b):

Let the numbers be x, x+1, x+2, x+3 therefore x+1+x+3=132 =>2x+4=132 =>2x=128 =>x=64

OpenStudy (2bornot2b):

So the numbers are 64, 65, 66, 67

OpenStudy (lalaly):

let the integer=x second = (x+1) fourth= (x+3) (x+1)+(x+3)=132 2x=132-4 2x=128 x=64 so the integers are 64,65,66,67

OpenStudy (anonymous):

hey lalaly!

OpenStudy (lalaly):

hey hannah:D:D

OpenStudy (anonymous):